# Limits on forming coreless terrestrial worlds in the TRAPPIST-1 system

### v1

@@ 4 Discussion
Whether or not FeO cores would actually form from silicate melts early in planet histories remains an open question for planet masses 3.5 M⊕. The current experimental constraints and thermodynamic model show the asymptotical trend of O partitioning towards a fully oxidised core (Fig. 1). This may allow for a coreless scenario (L. T. Elkins‐Tanton & S. Seager, 2008) at higher P-T conditions within more massive planets. To verify whether FeO can differentiate to form a core or not, new high-pressure experiments under expanded P–T–composition space would be necessary in the future (see Luo et al. in prep.). If verified, experimentally or theoretically, the predicted FeO core, existing when MP 3.5 M⊕(Fig. 2), could readily become miscible with silicate melts, potentially refuting the conventional assumption that interiors of large rocky planets have to differentiate into a layered structure.
@@
Similarly, based on mixing properties in the binary MgSiO3–Fe (in fact MgO–Fe), MgSiO3–H2, and Fe-H2 systems, E. D. Young et al. (2025) concluded that MgSiO3, Fe, and H2 become fully miscible in deep interiors of molten cores (anything beneath the outer envelope, i.e. including silicate mantles and metallic cores) of sub-Neptunes or their super-Earth descendants. Considering the enhanced oxygen partitioning at high P-T conditions, FeO instead of pure Fe, is a more likely end-member and may enable full miscibility at even shallower depths.
@@
The partitioning of oxygen into the metal phase has important implications for atmospheres of sub-Neptunes and super-Earths. Firstly, the oxidation of metallic iron limits the amount of O available to oxidize H. As a result, the amount of atmospheric that can be outgassed would decrease with increasing planet mass. This is broadly consistent with H. E. Schlichting & E. D. Young (2022), where water concentrations in silicate mantles and envelopes of super-Earths are significantly lower if one takes into account the chemical equilibration between the core, mantle, and atmosphere. Secondly, as noted by A. Werlen et al. (2025), parameters such as planet mass, relative mass fractions of core–mantle–atmosphere, all strongly influence the atmospheric C/O ratio. In a closed system with fixed initial amounts of C and O (L. Schaefer & B. Fegley, 2017), if the metal–silicate (metallic core–silicate magma ocean) equilibration is considered, the partitioning of both elements into iron would determine the amounts of C and O available in the magma ocean, which would eventually modify the observed atmospheric C/O ratios owing to magma ocean-atmosphere interactions (C. Seo et al., 2024; T. Lichtenberg & Y. Miguel, 2025; M. C. Nixon et al., 2025).
@@
A caveat with the current model of O partitioning is that it is established on the basis of experimental data on relatively ‘dry’ terrestrial materials (D. Huang & J. Badro, 2018; D. Huang et al., 2020, 2021, 2024), and is hence applicable to rocky interiors or lava planets in the regime of Henry’s law; its applicability to ocean planets (M. J. Kuchner, 2003; A. Léger et al., 2004), formed beyond the snow line and containing 10s wt% level of water, may be limited and requires further investigations by the mineral physics community in the future. As a step towards a more realistic scenario, future studies should address interactions among volatiles (H, C, N and O) when metal and silicate react with each other (e.g. D. Huang et al., 2024), and how they affect the partitioning and thereby atmospheric compositions.
5 Conclusions
@@
We present a new equilibration model for Fe–FeO core phases that predicts the oxygen content of the metal phase under given pressure–temperature conditions. The model is constrained by experimental data spanning pressures from a few GPa up to 104 GPa (D. Huang et al., 2024). Our results show a clear trend of increasing oxygen incorporation with increasing planetary mass, under the assumption that equilibration pressure scales with mass. Although extrapolation is required beyond 2 M⊕, where equilibration pressures exceed experimental limits, the model predicts that pure Fe cores become impossible above a critical planet mass. Specifically, the extrapolation suggests this threshold occurs near 3.5 M⊕, though the exact value depends on the assumptions adopted here. The key result is that beyond this critical mass, the metallic core will be saturated with oxygen.
@@
The implications for the TRAPPIST-1 planets are that coreless scenarios can be ruled out, and that the FeO content among the planets likely ranges from 0.01 to 0.05 wt%. Previously, the possible interior configurations included coreless models, variations in bulk Fe/Mg (or Fe/Si) ratios, and volatile-rich compositions. However, given recent spectroscopic observations from JWST and evolutionary constraints on water loss, we are left with a single plausible explanation: the Fe/Mg (or Fe/Si) ratios are sub-solar. This interpretation is consistent with previous estimates of the stellar composition (C. T. Unterborn et al., 2018b; C. Dorn et al., 2018b; L. Acuña et al., 2021). The fact that super-Earth metal phases become oxygen saturated for high mass planets has possible implications for any atmosphere that can degas from an early mantle magma ocean, including the observed atmospheric H2O and C/O ratio.

### v2

@@ IV Discussion
Whether or not FeO cores would actually form from silicate melts early in planet histories remains an open question for planet masses 3.5 M⊕. The current experimental constraints and thermodynamic model show the asymptotical trend of O partitioning towards a fully oxidised core (Fig. 1). It appears to be unlikely that there exists a coreless scenario at higher P-T conditions within more massive Earth-like planets, which is different from the water-rich scenario proposed by ElkinsTanton2008 . To verify whether FeO can differentiate to form a core or not, new high-pressure experiments under expanded P–T–composition space would be necessary in the future (see Luo et al. in prep.). If verified, experimentally or theoretically, the predicted FeO core, existing when MP 3.5 M⊕(Fig. 2), could readily become miscible with silicate melts, potentially refuting the conventional assumption that interiors of large rocky planets have to differentiate into a layered structure.
@@
Similarly, based on mixing properties in the binary MgSiO3–Fe (in fact MgO–Fe), MgSiO3–H2, and Fe-H2 systems, young2025 concluded that MgSiO3, Fe, and H2 become fully miscible in deep interiors of molten cores (anything beneath the outer envelope, i.e. including silicate mantles and metallic cores) of sub-Neptunes or their super-Earth descendants. Considering the enhanced oxygen partitioning at high P-T conditions, FeO instead of pure Fe, is a more likely end-member and may enable full miscibility at even shallower depths.
@@
The partitioning of oxygen into the metal phase has important implications for atmospheres of sub-Neptunes and super-Earths. Firstly, the oxidation of metallic iron limits the amount of O available to oxidize H. As a result, the amount of atmospheric that can be outgassed would decrease with increasing planet mass. This is broadly consistent with schlichting_chemical_2022, where water concentrations in silicate mantles and envelopes of super-Earths are significantly lower if one takes into account the chemical equilibration between the core, mantle, and atmosphere. Secondly, as noted by Werlen2025, parameters such as planet mass, relative mass fractions of core–mantle–atmosphere, all strongly influence the atmospheric C/O ratio. In a closed system with fixed initial amounts of C and O (schaefer_redox_2017), if the metal–silicate (metallic core–silicate magma ocean) equilibration is considered, the partitioning of both elements into iron would determine the amounts of C and O available in the magma ocean, which would eventually modify the observed atmospheric C/O ratios owing to magma ocean-atmosphere interactions (Seo2024; Lichtenberg2025; Nixon2025; Bower_2025).
@@
A caveat with the current model of O partitioning is that it is established on the basis of experimental data on relatively ‘dry’ terrestrial materials (Huang2018; Huang2020; Huang2021; Huang2024), and is hence applicable to rocky interiors or lava planets in the regime of Henry’s law; its applicability to ocean planets (Kuchner2003; Leger2004), formed beyond the snow line and containing 10s wt% level of water, may be limited and requires further investigations by the mineral physics community in the future. As a step towards a more realistic scenario, future studies should address interactions among volatiles (H, C, N and O) when metal and silicate react with each other (e.g. Huang2024), and how they affect the partitioning and thereby atmospheric compositions.
V Conclusions
@@
We present a new metal–silicate equilibration model that predicts the oxygen content of the metal phase under given pressure–temperature conditions. The model is constrained by experimental data spanning pressures from a few GPa up to 104 GPa (Huang2024). Our results show a clear trend of increasing oxygen incorporation with increasing planetary mass, under the assumption that equilibration pressure scales with mass. Although extrapolation is required beyond 2 M⊕, where equilibration pressures exceed experimental limits, the model predicts that pure Fe cores become impossible above a critical planet mass. Specifically, the extrapolation suggests this threshold occurs near 3.5 M⊕, though the exact value depends on the assumptions adopted here. The key result is that as the planet approaches this critical mass, its metallic core will become saturated with oxygen.
@@
The implications for the TRAPPIST-1 planets are that coreless scenarios can be ruled out, and that the oxygen content among their cores likely ranges from 0.01 to 0.05 wt%. Previously, the possible interior configurations included coreless models, variations in bulk Fe/Mg (or Fe/Si) ratios, and volatile-rich compositions. However, given recent spectroscopic observations from JWST and evolutionary constraints on water loss, we are left with a single plausible explanation: the Fe/Mg (or Fe/Si) ratios are sub-solar. This interpretation is consistent with previous estimates of the stellar composition (unterborn_inward_2018; dorn_interior_2018; acuna2021characterisation). The fact that super-Earth metal phases become oxygen saturated for high mass planets has possible implications for any atmosphere that can degas from an early mantle magma ocean, including the observed atmospheric H2O and C/O ratio.
All authors contributed equally to the manuscript.